OpenStudy (anonymous):

Find the center and radius of the circle whose equation is x 2 + y 2 - 6x - 2y + 4 = 0

OpenStudy (accessdenied):

So, you would want to put this into the vertex form. That would tell you the information you want by observation. Do you know how to do that?

OpenStudy (anonymous):

Just to comfirm, by x2 you mean \[x ^{2}\]? If that is what it means then this is what you do. First I would put all the terms with an x and all the terms with a y into little groups. \[(x ^{2} -6x) + (y ^{2} -2y) +4 = 0\] Then you do something called completing the square for both the brackets. So to complete the square we half the single x and y terms and take the squares out of the bracket. \[(x -3)^{2} + (y -1) ^{2} +4 =0\] However when we multiple out those brackets we will get a 9 and a 1 appearing that weren't there before. (-3x-3 =9) (-1x-1=1) So to counteract those terms we add a -9 and a -1 \[(x -3)^{2} + (y -1) ^{2} +4-1-9 =0\] Finally we just rearrange to get all the numbers on one side\[(x -3)^{2} + (y -1) ^{2} =6\] This is now in the form of the equation of a circle. So the centre is (3,1) and the radius is root 6.

OpenStudy (anonymous):

no i dont

OpenStudy (anonymous):

these are the answers i have to choose from. (3,1), r = sqrt 6 (-3,-1), r = sqrt 6 (3,1) r = 6 (9,1), r = 36

OpenStudy (accessdenied):

We would need a method called completing the square. We want to create perfect square trinomials out of the terms with x and the terms with y: x^2 - 6x, y^2 - 2y The method is fairly straightforward, we divide the coefficient multiplied to x by 2 and then square the result. That will need to be added and subtracted because we can factor the first part: \(\color{blue}{x^2 - 2h x + h^2} - h^2 = \color{blue}{(x - h)^2} -h^2 \) In the problem, we take -6x. Its coefficient is -6. We divide it by two, then square the result: -6 / 2 = -3; (-3)^2 = 9. The same process applies to the -2y term. So if we add and subtract 9 from the x-part, \(\color{blue}{x^2 - 6x + 9} - 9 \), we make a completed square: \( \color{blue}{(x - 3)^2} - 9\)

OpenStudy (accessdenied):

So to show that in context, \( x^2 + y^2 - 6x - 2y + 4 = 0 \) Group the x and y terms: \( x^2 - 6x + y^2 - 2y + 4 = 0 \) We add and subtract that which makes a perfect square, which was discussed to be 9. \( x² - 6x \color{lime}{+ 9 - 9} + y² - 2y + 4 = 0\) \( \color{blue}{(x^2 - 6x + 9)} - 9 + y^2 - 2y + 4 = 0 \) \( \color{blue}{(x - 3)^2} - 9 + y^2 - 2y + 4 = 0\) If you can do this same process for the y quadratic, you will be able to move your constants onto the right and finish!